# Cognitive Effort in the Two-Step Task: An Active Inference Drift-Diffusion Model Approach

### v1

@@ Abstract
High-level theories rooted in the Bayesian Brain Hypothesis often frame cognitive effort as the cost of resolving the conflict between habits and optimal policies. In parallel, evidence accumulator models (EAMs) provide a mechanistic account of how effort arises from competition between the subjective values of available options. Although EAMs have been combined with frameworks like Reinforcement Learning to bridge the gap between high-level theories and process-level mechanisms, relatively less attention has been paid to their implications for a unified notion of cognitive effort. Here, we combine Active Inference (AIF) with the Drift-Diffusion Model (DDM) to investigate whether the resulting AIF-DDM can simultaneously account for effort arising from both habit violation and value discriminability. To our knowledge, this is the first time AIF has been combined with an EAM. We tested the AIF-DDM on a behavioral dataset from the two-step task and compared its predictions to an information-theoretic definition of cognitive effort based on AIF. The model’s predictions successfully accounted for second-stage reaction times but failed to capture the dynamics of the first stage. We argue the latter discrepancy likely stems from the experimental design rather than a fundamental flaw in the model’s assumptions about cognitive effort. Accordingly, we propose several modifications of the two-step task to better measure and isolate cognitive effort. Finally, we found that integrating the DDM significantly improved parameter recovery, which could help future studies to obtain more reliable parameter estimates.
@@
active inference drift-diffusion model two-step task cognitive effort
@@ 1 Introduction
Building upon the Bayesian Brain Hypothesis (BBH), numerous studies in the past decade have sought to formalize cognitive effort using information-theoretic principles [18]. According to the BBH, humans maintain an internal world model encoded in prior beliefs, which they continuously update as they interact with the environment. Within this context, cognitive effort arises from the conflict between a pre-existing belief about how to act (a habit) and an updated belief about the most optimal policy [6].
@@
However, cognitive effort not only depends on the divergence between prior and posterior beliefs, but also on the competition between the subjective values of the available options. Specifically, effort increases as the difference between these subjective values—or value discriminability—decreases. From an information-theoretic perspective, this can be framed as the effort required to resolve uncertainty in the prior belief about which action to take [18]. Yet, information theory provides no explicit description of the underlying deliberation process, making it challenging to link its predictions to specific neural signatures. In contrast, evidence accumulator models (EAMs) offer a powerful mechanistic explanation for how value competition gives rise to cognitive effort that can be empirically tested (e.g., [13]).
@@
A major limitation of using EAMs to study cognitive effort is that they do not typically account for how beliefs are formed. This limitation is often addressed by combining EAMs with Reinforcement Learning (RL) [7]. However, in recent years, Active Inference (AIF) has emerged as a powerful alternative to RL, offering a first-principles account of perception, learning, and decision-making [4].
@@
Here, we investigate whether integrating a Drift Diffusion Model (DDM), a prominent class of EAM, with AIF can simultaneously describe the influence of both value discriminability and habit violation on cognitive effort. To our knowledge, this is the first time AIF has been combined with an EAM to model human behaviour. We test the combined AIF-DDM using the two-step task, a version of the multi-armed bandit where participants have to plan two steps ahead [2]. Furthermore, we compare its predictions with a recently developed definition of cognitive effort in AIF [6].
@@
Our work builds directly on a recent study that developed an AIF model of the two-step task [3]. The study demonstrated that AIF outperformed a Hybrid Reinforcement Learning (HRL) model (which combines model-free and model-based strategies [11]) in two out of the four tested datasets, while achieving comparable performance in the remaining two. Furthermore, the authors presented compelling evidence for directed exploration—a key differentiator between AIF and HRL. Despite these achievements, the study reported that it was not possible to identify which specific AIF learning mechanisms participants were using. Therefore, we tested whether DDM integration can help to solve this ambiguity. This additional goal was motivated by prior research showing that integrating a DDM with HRL improves the reliability of model-based estimates, due to the additional information provided by Reaction Times (RTs) [11].
@@ 2.1 Participants and behavioural task
As the name suggests, each trial of the two-step task comprises two stages (Figure 1). In the first stage, participants choose between two actions. Each action leads to one of two second-stage states via a probabilistic transition, which is either common ($p=0.7$) or rare ($p=0.3$). Importantly, the two first-stage actions have opposite most-likely transitions. These transition probabilities remain constant throughout the task. After transitioning to a second-stage state, participants make a final choice between two actions. Each of these second-stage actions results in a monetary reward or no reward, depending on its current outcome probability. In contrast to the fixed transitions, these outcome probabilities fluctuate independently over time, following Gaussian random walks.
@@
We analysed data from the "Magic Carpet" dataset, which was made available by [12] and originally comprised 24 participants. In this experiment, participants had a 2-second deadline to respond at each stage. Every time this deadline was surpassed, the trial was labelled as a missed trial and the participant moved on to the next stage or trial. Therefore, before conducting any analysis, we identified and removed these trials. We also considered invalid trials those with at least one RT smaller than 100ms (to exclude anticipatory responses that are too fast to reflect genuine deliberation [5]). Finally, any participant with more than 10% invalid trials was removed from the dataset. One participant was excluded under this criterion, with 44.1% of their trials being either missed or too fast. Therefore, the original sample size was reduced to n=23. In total, 6.53% of the data was excluded from the analysis (including all the data from the removed participant).
2.2 Active inference drift-diffusion model of the two-step task
@@
In what follows, we will use the same notation as in [3]. For a given trial $t$, the first-stage state and chosen actions are denoted by $s_{1,t}$ and $a_{1,t}$, respectively. Likewise, the second-stage state and chosen action are denoted by $s_{2,t}$ and $a_{2,t}$, respectively. Note that the first-stage state can only be $s_{\text{A}}$, but the second-stage state may be $s_{\text{B}}$ or $s_{\text{C}}$, depending on the transition (see Figure 1). Outcomes $o_{t}\in\{0,1\}$ represent the observed reward (or absence of it) after choosing a second-stage action.
@@
Where $\mathcal{A}_{B}$ and $\mathcal{A}_{C}$ are the sets of available actions in the second-stage states $s_{B}$ and $s_{C}$ respectively, and the second-stage EFEs, $G(a_{2})$ are given by Eq. 1.
Previous studies have recorded a tendency for participants to repeat initial-stage actions independent of the outcome [3]. In AIF, this behaviour is captured by a habit term, E, which in the model developed by [3] is governed by a free parameter, $\kappa$, which controls the strength of an agent’s reliance on habits (see Eq. 6, Appendix 0.A).
At the first stage, the agent will select an action that minimizes the EFE corrected by the habitual bias $G_{t}^{\text{net}}(s_{1},a_{1})=G_{t}(s_{1},a_{1})+E(a_{1})$. At the second stage, it is assumed that there is no habitual bias, and therefore $G_{t}^{\text{net}}(s_{2},a_{2})=G_{t}(s_{2},a_{2})$.
@@ 2.3 Cognitive effort and active inference
For the first stage of the two-step task, the AIF-DDM aligns with the first prediction of Eq. 4. This is because for a choice between actions $a_{A}$ and $a_{B}$, with $\mathbf{G}=[G_{A},G_{B}]$, and $\mathbf{E}=[E_{A},E_{B}]$, RTs should increase as the magnitude of the drift rate decreases, which is proportional to $|\Delta G^{\text{net}}|=|G_{A}+E_{A}-G_{B}-E_{B}|$. Therefore, higher RTs may occur when, for example $|G_{A}|\gg|G_{B}|$, $|E_{A}|\ll|E_{B}|$ and $|G_{A}|\approx|E_{B}|$. In other words, when there is an incongruence between G and E.
However, our model contradicts the second prediction of Eq. 4, because for a constant E, RTs should increase as $|G_{A}|$ gets closer to $|G_{B}|$. Interestingly, for the second stage, Eq. 4 predicts minimal effort if we assume no habits (i.e. $E_{A}=E_{B}=0$), in contrast to AIF-DDM, where effort can still be high if the context-senstive EFEs are closely matched.
@@ 2.4 Model fitting and comparison procedures
We followed the same Maximum Likelihood Estimate (MLE) procedure as in [3], to fit the models’ free parameters to the behavioural data. For each participant and model, we found the parameter set with the highest likelihood using Scipy’s ’L-BFGS-B’ algorithm. This step was repeated 35 times for each participant, with different (uniformly) randomized initializations for all parameters. After completing all the runs, we selected the parameter set with the maximum likelihood and used its value for model comparison.
For the pure AIF and HRL models, we computed the likelihoods using the choice probability distributions of each model. For AIF-DDM and HRL-DDM however, the likelihoods were given by the Wiener’s First-Passage Time Distribution (WFPT) provided by the HDDM Python package [16]. Further details can be found in the Appendix, section 0.C.
For model comparison, we relied on the Statistical Parametric Mapping (SPM) [15] and Variational Bayesian Analysis (VBA) [1] toolboxes, with MATLAB R2022b. We performed a group-level random-effect Bayesian model selection (BMS) procedure, which requires the log-model evidence (LME) of each model for each participant. We used two LME approximations: the Bayesian Information Criterion (BIC) and the Akaike’s Information Criterion (AIC) scores [8] (see Appendix, section 0.C). Using the BMS procedure, for each model we calculated Expected Posterior Probability (EPP), Estimated Model Frequency (EMF)- the proportion of participants for whom the model was the best-fitting, Exceedance Probability (EP), which measures how likely it is that the model is more frequent than all the other considered across all participants and Protected Exceedance Probabilities (PEP), similar to EP but accounting for the possibility that apparent differences in model frequencies arise due to chance [9].
@@
2.5 Model and Parameter recovery procedure
@@
We fitted all the models to each model’s simulated dataset for the model recovery analysis. For each simulated dataset, we then calculated the fraction of times each of the models was the best-fitting, and plotted the results in a confusion matrix (Figure 2).
@@ 3.1 Model recovery analysis
Model recovery analysis indicates that $\text{AIF-DDM}_{\text{FULL}}$ is frequently misclassified as either $\text{AIF-DDM}_{\text{NSD}}$ or $\text{AIF-DDM}_{\text{NSD}}$ (Figure 2). This poor recovery likely stems from the full model encompassing all update rules present in the other two models [3]. The rest of the models had a better but modest recovery (except HRL-DDM, which had a perfect recovery). Thus, combining AIF with a DDM could not help substantially to differentiate between different learning variants as we hoped for. Overall, both AIC and BIC scores show similar results. Although the AIC score seems slightly more reliable since it could achieve a better recovery for both $\text{AIF-DDM}_{\text{FULL}}$ and $\text{AIF-DDM}_{\text{NUD}}$.
Figure 2: Model recovery results. Each cell contains the fraction of participants (simulated under a given true model) for which a given model was classified as the best-fitting, according to the BIC score (left) or the AIC score (right).
@@ 3.2 Model comparison results
The BMS analysis revealed different results depending on the score used. According to the BIC score, $\text{AIF-DDM}_{\text{NSD}}$ and $\text{AIF-DDM}_{\text{NPS}}$ dominated moderately across participants with Expected Posterior Probabilities, $\text{EPP}=[0.39,0.31]$, Protected Exceedance Probabilities, $\text{PEP}=[0.63,0.29]$ and Estimated Model Frequencies, $\text{EMF}=[0.44,0.35]$, respectively (see Figure 5, first row). However, according to the AIC scores $\text{AIF-DDM}_{\text{FULL}}$ was the best-fitting model with $\text{EPP}=0.57$, $\text{PEP}=0.97$ and $\text{EMF}=0.7$ (see Figure 5, second row). The disagreement between AIC and BIC scores was also reported for the model comparison between pure HRL and AIF models [3], and likely stems from the fact that BIC penalizes complexity more than AIC.
Although the two BMS procedures yielded conflicting results, we selected $\text{AIF-DDM}_{\text{FULL}}$ as the best-fitting model, and focus exclusively on its results in the subsequent sections. This decision was based on the model recovery analysis, which showed that the full model is frequently misclassified as either $\text{AIF-DDM}_{\text{NSD}}$ or $\text{AIF-DDM}_{\text{NPS}}$. Nonetheless, we acknowledge that the BIC-based comparison favoured $\text{AIF-DDM}_{\text{NSD}}$ and $\text{AIF-DDM}_{\text{NPS}}$, so we replicated our analysis for these models and show their results in the Appendix. We excluded $\text{AIF-DDM}_{\text{NUD}}$ from further analysis as it performed poorly in both BMS procedures. Finally, the model family selection analysis revealed similar results to those reported in the study that compared pure AIF and HRL [3]. We found that the AIF-DDM family outperformed the HRL-DDM model according to both BIC (EP = 1, EMF = 0.9, EPP = 0.88) and AIC (EP = 1, EMF = 0.92, EPP = 0.9) scores.

### v2

@@ Abstract
High-level theories rooted in the Bayesian Brain Hypothesis often frame cognitive effort as the cost of resolving the conflict between habits and optimal policies. In parallel, evidence accumulator models (EAMs) provide a mechanistic account of how effort arises from competition between the subjective values of available options. Although EAMs have been combined with frameworks like Reinforcement Learning to bridge the gap between high-level theories and process-level mechanisms, relatively less attention has been paid to their implications for a unified notion of cognitive effort. Here, we combine Active Inference (AIF) with the Drift-Diffusion Model (DDM) to investigate whether the resulting AIF-DDM can simultaneously capture the effort arising from both habit violation and value discriminability. To our knowledge, this is the first time AIF has been combined with an EAM. We tested the AIF-DDM on a behavioral dataset from the two-step task and compared its predictions to an information-theoretic definition of cognitive effort based on AIF. The model’s predictions successfully accounted for second-stage reaction times but failed to capture the dynamics of the first stage. We argue the latter discrepancy likely stems from the experimental design rather than a fundamental flaw in the model’s assumptions about cognitive effort. Accordingly, we propose several modifications of the two-step task to better measure and isolate cognitive effort. Finally, we found that integrating the DDM significantly improved parameter recovery, which could help future studies to obtain more reliable parameter estimates.
@@
Active inference Drift-diffusion model Cognitive effort
@@ 1 Introduction
Building upon the Bayesian Brain Hypothesis (BBH), numerous studies in the past decade have tried to formalize cognitive effort using information-theoretic principles [25]. According to the BBH, humans maintain an internal world model encoded in prior beliefs, which they continuously update as they interact with the environment. Within this context, cognitive effort arises from the conflict between a pre-existing belief about how to act (a habit) and an updated belief about the optimal policy [10].
In a decision-making task, cognitive effort may also arise from the competition between the subjective values of the available choices. When these values are closer together—that is, when value discriminability is low—reaction times (RTs) tend to increase (e.g., [7, 3]). Although RTs are not a direct measure of cognitive effort, they are often used as a proxy (e.g., [24]), based on the assumption that slower responses reflect more information processing.
@@
From an information-theoretic perspective, the effect of value discriminability on RTs can be understood as the additional cognitive effort required to resolve increased choice uncertainty [13, 8, 6]. Yet, information theory provides no explicit account of the underlying deliberation process, which complicates the task of linking its predictions to specific neural signatures. In contrast, evidence accumulator models (EAMs) offer a mechanistic explanation for how value competition shapes decision speed, which may be empirically tested (e.g., [18]).
@@
A major limitation of using EAMs to study cognitive effort is that they are agnostic to how beliefs are formed. This limitation is often addressed by combining EAMs with Reinforcement Learning (RL) [11]. However, in recent years, Active Inference (AIF) has emerged as a powerful alternative to RL, offering a first-principles perspective on perception, learning, and decision-making [5].
@@
Here, we investigate whether integrating a Drift Diffusion Model (DDM), a prominent class of EAM, with AIF can simultaneously capture the influence of both value discriminability and habit violation on cognitive effort. To our knowledge, this is the first attempt to combine AIF with an EAM to model human behaviour. We evaluate the integrated AIF-DDM using the two-step task, a version of the multi-armed bandit in which participants must plan two steps ahead [2]. Furthermore, we compare its predictions with a recently proposed definition of cognitive effort in AIF [10].
@@
Our work builds directly on a recent study that developed an AIF model of the two-step task [4]. The study demonstrated that AIF outperformed a Hybrid Reinforcement Learning (HRL) model (which combines model-free and model-based strategies [16]) in two out of four datasets, while achieving comparable performance in the remaining two. Moreover, the authors provided compelling evidence for directed exploration—a key differentiator between AIF and HRL. Despite these achievements, the study could not determine which specific AIF learning mechanisms participants were using. Given prior evidence that integrating a DDM into an HRL model improved the reliability of model-based estimates [16], we tested whether the combined AIF–DDM could resolve this ambiguity.
@@ 2.1 Participants and behavioural task
As the name suggests, each trial of the two-step task consists of two stages (Fig. 1). In the first stage, participants choose between two actions. Each action leads to one of the two second-stage states through a probabilistic transition that is either common ($p=0.7$) or rare ($p=0.3$). Importantly, the two first-stage actions have opposite most-likely transitions. These transition probabilities remain constant throughout the task. After transitioning to a second-stage state, participants make a final choice between two actions. Each of these second-stage actions results in a monetary reward or no reward, depending on its current outcome probability. In contrast to the fixed transitions, these outcome probabilities fluctuate independently over time, following Gaussian random walks.
@@
We analysed data from the "Magic Carpet" dataset, which was made available by [17] and originally comprised 24 participants. In this experiment, participants had a 2-second deadline to respond at each stage. Every time this deadline was surpassed, the trial was labelled as a missed trial and the participant moved on to the next stage or trial. Before conducting any analysis, we identified and removed these trials. We also considered invalid trials those with at least one RT smaller than 100 ms (to exclude anticipatory responses that are too fast to reflect genuine deliberation [9]). Finally, any participant with more than 10% invalid trials was removed from the dataset. One participant was excluded under this criterion, with 44.1% of their trials being either missed or too fast. Therefore, the original sample size was reduced to n=23. In total, 6.53% of the data was excluded from the analysis (including all the data from the removed participant).
2.2 An active inference drift-diffusion model of the two-step task
@@
In what follows, we will use the same notation as in [4]. For a given trial $t$, the first-stage state and chosen actions are denoted by $s_{1,t}$ and $a_{1,t}$, respectively. Likewise, the second-stage state and chosen action are denoted by $s_{2,t}$ and $a_{2,t}$, respectively. Note that the first-stage state can only be $s_{\text{A}}$, but the second-stage state may be $s_{\text{B}}$ or $s_{\text{C}}$, depending on the transition (see Fig. 1). Outcomes $o_{t}\in\{0,1\}$ represent the observed reward (or absence of it) after choosing a second-stage action.
@@
Where $\mathcal{A}_{B}$ and $\mathcal{A}_{C}$ are the sets of available actions in the second-stage states $s_{B}$ and $s_{C}$ respectively, and $G(a_{2})$ are the second-stage EFEs given by Eq. 1.
Previous studies have reported a tendency for participants to repeat initial-stage actions, regardless of the outcome history [4]. Within AIF, this behaviour is formalised through a habit term $E$. In the model of [4], $E$ is parameterised by $\kappa$, which modulates the extent of an agent’s reliance on habits (see Eq. 6, Appendix 0.A).
At the first stage, an agent will select an action that minimizes the EFE corrected by the habitual bias $G_{t}^{\text{net}}(s_{1},a_{1})=G_{t}(s_{1},a_{1})+E(a_{1})$. At the second stage, no habitual bias is assumed, and therefore $G_{t}^{\text{net}}(s_{2},a_{2})=G_{t}(s_{2},a_{2})$.
@@ 2.3 Cognitive effort and active inference
For the first stage of the two-step task, the AIF-DDM aligns with the first prediction of Eq. 4. For a choice between actions $a_{A}$ and $a_{B}$, with $\mathbf{G}=[G_{A},G_{B}]$, and $\mathbf{E}=[E_{A},E_{B}]$, RTs should increase as the magnitude of the drift rate decreases, which is proportional to $|\Delta G^{\text{net}}|=|G_{A}+E_{A}-G_{B}-E_{B}|$. Therefore, higher RTs may occur when, for example $|G_{A}|\gg|G_{B}|$, $|E_{A}|\ll|E_{B}|$ and $|G_{A}|\approx|E_{B}|$. In other words, when there is an incongruence between G and E.
However, our model contradicts the second prediction of Eq. 4, because for a constant E, RTs should increase as $|G_{A}|$ gets closer to $|G_{B}|$. Interestingly, for the second stage, Eq. 4 predicts minimal effort if we assume no habits (i.e. $E_{A}=E_{B}=0$), in contrast to AIF-DDM, where effort can still be high if the context-sensitive EFEs are closely matched.
@@ 2.4 Model fitting and comparison procedures
We followed the same Maximum Likelihood Estimate (MLE) procedure as in [4], to fit the models’ free parameters to the behavioural data  [^1] Code available at https://github.com/decide-ugent/aif-ddm. For each participant and model, we found the parameter set with the highest likelihood using Scipy’s ’L-BFGS-B’ algorithm [19]. This step was repeated 35 times for each participant, with different (uniformly) randomized initializations for all parameters. After completing all the runs, we selected the parameter set with the maximum likelihood and used its value for model comparison.
For the pure AIF and HRL models, we computed the likelihoods using the choice probability distributions of each model. For AIF-DDM and HRL-DDM however, the likelihoods were given by the Wiener’s First-Passage Time Distribution (WFPT) provided by the HDDM Python package [22]. Further details can be found in Appendix 0.C.
For model comparison, we relied on the Statistical Parametric Mapping (SPM) [21] and Variational Bayesian Analysis (VBA) [1] toolboxes, using MATLAB R2022b. We performed a group-level random-effect Bayesian model selection (BMS) procedure, which requires the log-model evidence (LME) of each model for each participant. We used two LME approximations: the Bayesian Information Criterion (BIC) and the Akaike’s Information Criterion (AIC) scores [12] (see Appendix 0.C). Using the BMS procedure, we estimated four model comparison metrics: the Expected Posterior Probability (EPP); the Estimated Model Frequency (EMF), or the proportion of participants best fit by the model; the Exceedance Probability (EP), which quantifies the likelihood that the model is more frequent than competing models across participants; and the Protected Exceedance Probability (PEP), similar to EP but accounting for the possibility that apparent differences in model frequencies arise due to chance [14].
@@
2.5 Model and parameter recovery procedures
@@
For the model recovery analysis, each model was first used to simulate a full experimental dataset with the same parameter sets as in the parameter recovery analysis. We then fitted all models to each simulated dataset and calculated the proportion of times each model was identified as best-fitting. Results were summarized in a confusion matrix (Fig. 2).
@@ 3.1 Model recovery analysis
Model recovery analysis indicates that $\text{AIF-DDM}_{\text{FULL}}$ is frequently misclassified as either $\text{AIF-DDM}_{\text{NSD}}$ or $\text{AIF-DDM}_{\text{NSD}}$ (Fig. 2). The poor recovery likely stems from the full model encompassing all update rules present in the other two models [4]. The rest of the models had a better but modest recovery (except HRL-DDM, which had a perfect recovery). Thus, combining AIF with a DDM could not help substantially to differentiate between different learning variants as we hypothesised. Overall, both AIC and BIC scores show similar results. Although the AIC score seems slightly more reliable since it could achieve a better recovery for both $\text{AIF-DDM}_{\text{FULL}}$ and $\text{AIF-DDM}_{\text{NUD}}$.
Figure 2: Model recovery results. Each cell contains the fraction of simulated experiments from a given true model, for which a corresponding model was classified as the best-fitting, according to the BIC score (left) or the AIC score (right).
@@ 3.2 Model comparison results
The BMS analysis revealed different results depending on the score used. According to the BIC score, $\text{AIF-DDM}_{\text{NSD}}$ and $\text{AIF-DDM}_{\text{NPS}}$ dominated moderately across participants with Expected Posterior Probabilities, $\text{EPP}=[0.39,0.31]$, Protected Exceedance Probabilities, $\text{PEP}=[0.63,0.29]$ and Estimated Model Frequencies, $\text{EMF}=[0.44,0.35]$, respectively (see Fig. 5, first row). However, according to the AIC scores $\text{AIF-DDM}_{\text{FULL}}$ was the best-fitting model with $\text{EPP}=0.57$, $\text{PEP}=0.97$ and $\text{EMF}=0.7$ (see Fig. 5, second row). The disagreement between AIC and BIC scores was also reported for the model comparison between pure HRL and AIF models [4], and likely stems from the fact that BIC penalizes complexity more than AIC.
Even though the two BMS procedures produced conflicting results, we selected $\text{AIF-DDM}_{\text{FULL}}$ as the best-fitting model, and focus exclusively on its results in the subsequent sections. This decision was based on the model recovery analysis, which showed that the full model is frequently misclassified as either $\text{AIF-DDM}_{\text{NSD}}$ or $\text{AIF-DDM}_{\text{NPS}}$. Nonetheless, we acknowledge that the BIC-based comparison favoured $\text{AIF-DDM}_{\text{NSD}}$ and $\text{AIF-DDM}_{\text{NPS}}$, so we replicated our analysis for these models and show their results in Appendix 0.D. We excluded $\text{AIF-DDM}_{\text{NUD}}$ from further analysis as it performed poorly in both BMS procedures. Finally, the model family selection analysis revealed similar results to those reported in the study that compared pure AIF and HRL [4]. We found that the AIF-DDM family outperformed the HRL-DDM model according to both BIC (EP = 1, EMF = 0.9, EPP = 0.88) and AIC (EP = 1, EMF = 0.92, EPP = 0.9) scores.